You have 150ml of pure water, add 5% of salt to the water. how many ml salt water will you have?
Answer by dkppathak(439)

total pure water 150 ml
5% salt added =5/100 of 150 ml
=7.5 ml
salted water solution will be =150+7.50 =157.50 ml
answer 157.50 ml